OpenStudy (anonymous):

HELP! How do I convert from intercept form, not slope intercept form, to general equation form in a question where the X intercept= 2,0 and Y Intercept= 0,3? Explanation too please?

OpenStudy (cwrw238):



OpenStudy (anonymous):

?

OpenStudy (anonymous):

Its not a graphing question, its a formula conversion

OpenStudy (cwrw238):

thats the graph of the line the slope = -3/2 we may write it as y = (-3/2)x + 3 +3 because thats the y intercept now to convert to standard form multiply through by 2:- 2y = -3x + 6 3x + 2y - 6 = 0

OpenStudy (anonymous):

how come the +3 also get multiplied along with y?

OpenStudy (cwrw238):

in order to keep the balance of the equation you must multiply each term by 2

OpenStudy (anonymous):

OH, I understand, thanks for the help!

OpenStudy (cwrw238):

i've converted the slope intercept form - not sure what you mean by intercept form.

OpenStudy (cwrw238):

yw

OpenStudy (mathmath333):

\(\large\rm \dfrac{x}{3}+\dfrac{y}{2}=1 \) \(\large\rm 2x+3y=6 \)

OpenStudy (mathmath333):

\(\large\rm \dfrac{x}{a}+\dfrac{y}{b}=1 \) a and b are x and y intercept

OpenStudy (cwrw238):

oh ok - that first equation is the intercept form - right? never come across that term beforre

OpenStudy (cwrw238):

dont seem to teach that in UK schools!

OpenStudy (anonymous):

I think both of you get the right answer but @mathmath333 has the correct form

OpenStudy (cwrw238):

yea - i've learned something new today

OpenStudy (mathmath333):

lol

OpenStudy (anonymous):

But I'm not sure where the six comes in from in @mathmath333 's equation

OpenStudy (cwrw238):

he's multiplied each term by the lcd OF 3 AND 2 TO GET RID OF THE PARENTHESES and get the standard form

OpenStudy (mathmath333):

multilpy both sides by 6

OpenStudy (mathmath333):

yes lcm